OpenStudy (anonymous):

using the matrix tool to solve the system of equation choose the correct ordered pair . 7x+2y=46 8x+8y=64 [ 7, 2,46 8, 8, 64 ] a ( 4,9) b ( 4,4) c ( 2,6) d ( 6,2)

OpenStudy (ybarrap):

$$ A=\begin{bmatrix} 7& 2 \\ 8& 8 \\ \end{bmatrix}\\ B= \begin{bmatrix} 46 \\ 64 \\ \end{bmatrix}\\ X=\text{ solution}:\\ AX=B\\ X=A^{-1}B $$ Where \(A^{-1}\) is the inverse of A. Using Wolfram|Alpha as the tool: http://www.wolframalpha.com/input/?i=matrix+multiplication&a=*C.matrix+multiplication-_*Calculator.dflt-&f2=inverse%7B%7B7%2C2%7D%2C%7B8%2C8%7D%7D&f=MatricesOperations.theMatrix1%5Cu005finverse%7B%7B7%2C2%7D%2C%7B8%2C8%7D%7D&f3=transpose%7B%7B46%2C64%7D%7D&f=MatricesOperations.theMatrix2_transpose%7B%7B46%2C64%7D%7D&a=*FVarOpt.1-_**-.***MatricesOperations.theMatrix3---.*-- Gives us as a solution: $$ X=\begin{bmatrix} 6 \\ 2 \\ \end{bmatrix} $$ Check our answer: $$ 7(6)+2(2)=46 ~\checkmark \\ 8(6)+8(2)=64 ~\checkmark $$ Make sense?

OpenStudy (anonymous):

that went way over my head lol

OpenStudy (ybarrap):

You need to use matrix tool to solve, correct?

OpenStudy (ybarrap):

Did you need to solve via some other method?

OpenStudy (anonymous):

no its matrix. it just looks different the way you explained it, idk maybe its just me